OpenStudy (anonymous):

@TechnoSoul Divide. (3a^3 + 6a^2) ÷ 3a A. a^2 + 6a^2 B. 3a^3 + 6a^2 – 3a C. a^2 + 2a D. 3a^3 + 6a^2 + 3a

OpenStudy (ipwnbunnies):

\[\frac{3a^3 + 6a^2}{3a}\] To make the division easier, since there's one term in the denominator, you can split this expression into two, and simplify. \[\frac{3a^3}{3a} + \frac{6a^2}{3a}\]

OpenStudy (ipwnbunnies):

When dividing variables, first, divide the coefficients, or the numbers out front. When dividing variables raised to a power, follow this rule of powers. \[\frac{x^{a}}{x^{b}} = x^{a-b}\]

OpenStudy (ipwnbunnies):

For example, x^5/x^3 = x^(5-3) = x^2

OpenStudy (anonymous):

so that would mean it's C?

OpenStudy (ipwnbunnies):

Yesh. c:

OpenStudy (anonymous):

Thnaks (: